# Search and Society Reimagining Information Access for Radical Futures

### v1

@@ Abstract
Information retrieval (IR) technologies and research are undergoing transformative changes. It is our perspective that the community should accept this opportunity to re-center our research agendas on societal needs while dismantling the artificial separation between the work on fairness, accountability, transparency, and ethics in IR and the rest of IR research. Instead of adopting a reactionary strategy of trying to mitigate potential social harms from emerging technologies, the community should aim to proactively set the research agenda for the kinds of systems we should build inspired by diverse explicitly stated sociotechnical imaginaries. The sociotechnical imaginaries that underpin the design and development of information access technologies needs to be explicitly articulated, and we need to develop theories of change in context of these diverse perspectives. Our guiding future imaginaries must be informed by other academic fields, such as democratic theory and critical theory, and should be co-developed with social science scholars, legal scholars, civil rights and social justice activists, and artists, among others. In this perspective paper, we motivate why the community must consider this radical shift in how we do research and what we work on, and sketch a path forward towards this transformation.
@@ 1 Introduction
Information retrieval (IR) research needs to explicitly articulate the sociotechnical imaginaries [191, 192] that underpin the design and development of information access technologies, and develop theories of change [377, 64, 343, 382] in context of these diverse perspectives. Jasanoff and Kim [192] define sociotechnical imaginaries as “collectively held, institutionally stabilized, and publicly performed visions of desirable futures, animated by shared understandings of forms of social life and social order attainable through, and supportive of, advances in science and technology”. These shared visions do not only imagine but also co-produce our futures through development and government of digital technologies [242]. Diverse imaginaries promoted by different corporations, professional communities, political organizations, and social movements can coexist “in tension or in a productive dialectical relationship” [192].
@@
In technological research and development, these diverse, and often diverging, perspectives and visions that guide the community frequently remain implicit and unstated [383] despite the significant influence they exert on what the community focuses on and produces. Because of the consequential role that access to information plays in political participation by citizens in democratic societies and social transformation [174, 292, 152, 83, 155], and as a social determinant of health [267] and economic progress [398, 271], it is even more important to critically reflect on the goals and motivations that guide the design and deployment of popular information access systems. What sociotechnical futures do IR researchers and system designers envision and how do they influence the design of current and future IR systems? Whose sociotechnical imaginaries are granted normative status and what myriad of radically alternative futures are we overlooking? For example, what are the implications of the reliance of popular search and social media platforms on advertising as the primary source of revenue generation [20] and how does big tech’s [279] increasing dominance over academic research [379] influences and/or homogenizes the kinds of IR systems we build? What is our role, as IR researchers, to safeguard communities from falling victim to crisis of imaginations [167] and how do we become more open to welcoming influences from radically new sociotechnical imaginaries? For example, what would IR systems look like if designed for futures informed by feminist, queer, decolonial, anti-racist, anti-casteist, and abolitionist thoughts, and if the focus of IR research was not to prop up colonial cisheteropatriarchal capitalist structures but to dismantle them?
@@
The Wave in the Mind: Talks and Essays on the Writer, the Reader, and the Imagination [222]
@@
For IR to concretely support diverse sociotechnical imaginaries the research community also needs to develop their own theories of change. Theory of change [377, 64, 343, 382] can be defined as a participatory process whereby groups and stakeholders articulate their long-term goals and identify necessary preconditions in a planning process. Consequently, in IR adjacent fields—e.g., in the Fairness, Accountability, and Transparency (FAccT) [^1] https://facctconference.org community—there has been similar recent calls to develop theories of change [383] to make explicit the diverse perspectives that motivate, and manifest in, the research undertaken by the community, as well as the visions for desired futures of responsible computing and the strategic pathways that lead to those desired futures. Here, we argue that IR research similarly needs to both explicitly articulate and support diverse imagined futures and develop corresponding theories of change for how new information access technologies can take us towards these desired worlds. Theories of change in this context may benefit IR research by encouraging community members to make their goals and assumptions explicit, making it feasible to test stated theories, encouraging the community to work towards building consensus, and even aid in developing potential means of evaluation of desired progress [377].
The idea that IR research must understand and contend with the social implications of the technology it produces is not new. Nearly half a century ago, Belkin and Robertson [40] concluded that IR research should acknowledge its responsibility to society and “must become both theoretically self-conscious and self-consciously based upon a social ideology”. More recently, at the third Strategic Workshop in Information Retrieval in Lorne (SWIRL) [91], researchers in attendance identified fairness, accountability, confidentiality, and transparency in IR (FACT IR) as socially consequential and strategically important research directions for the field. A subsequent workshop (FACTS-IR [276]) further emphasized the importance of developing actionable research agendas in these directions, adding “safety” as a fifth pillar. Indeed, a large body of recent IR literature grapples with questions of fairness [119, 402, 403, 404, 289, 103, 1, 375, 228, 97], as well as transparency and explainability [406, 18], in the context of IR. The increasing focus on these sociotechnical aspects of information access has been at least partly in response to recent advances in foundation models [61] and their implications for the future of information access [317].
It is our perspective that this growing focus on fairness and ethics in IR—despite having played a critical role in bringing much-needed attention to the societal implications of IR systems, and advancing the conversation about the IR community’s responsibility to broader society—operates within a severely constrained frame that leaves many underlying assumptions about the sociopolitical and economic incentives that guide IR research largely unchallenged. In machine learning, there has been similar recent perspectives [59, 258] that, for example, calls for shifting the lens from fairness and bias to the power differentials that exists between those who build technology, those who use it, and those who are subject to it. Others in the machine learning community have brought attention to the questions of how these technologies shift power [200], and simultaneously constrains ethics interventions in practice [381] and shapes our collective sociotechnical futures. Even by accepting the frame that we should develop fairness and transparency mechanisms for certain systems, we may inadvertently ignore the alternative perspective that some of these technologies should be dismantled, not made fairer, nor more transparent [38, 383, 256].
@@
IR research, instead of adopting a reactionary strategy of trying to mitigate potential social harms from emerging technologies by developing new fairness and transparency interventions, should aim to proactively set the research agenda for the kinds of systems we should build inspired by diverse explicitly stated sociotechnical imaginaries. Maximizing social good and minimizing harm in this context should not just be the concerns of the few in our community working on fairness, transparency, ethics, and related areas, but the domain of all IR research that should be guided by theories of change towards these envisioned futures. In this context, we largely agree with the perspective of Belkin and Robertson [40] but diverge on a critical point which is that we believe IR research should not seek any misguided notions of singular “universal” social ideology but explictly adopt pluralistic values and make space for diverse visions and perspectives.
@@
Consequently, the task for IR researchers here is not to put themselves in positions to pick the guiding social ideologies nor push technosolutionism to address today’s social problems. Instead, our guiding future imaginaries must be co-developed with social science scholars, legal scholars, civil rights and social justice activists, and artists, to name a few. Not all sociotechnical imaginaries are equal in this respect, and what futures guide our research must be informed by the values and ethics of our community, which should be constantly discussed, debated, and challenged by the community as part of our sociotechnical research and be open to external critique. To summarize, as a research community we should invest our energies and resources to: (i) Nurture digital spaces where radical visions and projects for human emancipation, social progress, and equity and justice can take shape, (ii) encourage experimentation within our research community with new approaches to information access informed by new sociotechnical imaginaries cross-pollinating through interdisciplinary scholarship, and (iii) ensure that the tools and artefacts we produce as a community do not aid systems of oppression nor cause other systemic social harms.
@@
Our contributions can be summarized as follows: We argue that IR research should explicate the sociotechnical futures we want to realize and develop theories of change towards these desired futures; we motivate why now is an appropriate time for the community to consider this shift; we draw from similar movements in IR-adjacent fields, specifically those with explicit values and prefigurative politics; we share some ideas on how the community can get started on this journey; and finally, who should be doing this work. Our goal with this paper is to raise sociopolitical consciousness in the IR research community so that we all see our research embedded in projects of future world making and recognize our collective responsibility to affect social good; and to dismantle the artificial separation between the work on fairness and ethics in IR and the rest of IR research.
@@ 2 Background
IR systems act as intermediaries between information seekers and information artefacts. These artefacts may represent: (i) economic and other opportunities for consumers, (ii) monetization opportunities for content creators and publishers, (iii) specific sociopolitical frames and viewpoints to consumers, and (iv) lenses to view individuals and groups that are subjects of the content. These systems infer the information need from highly incomplete expressions of interests (e.g., short keyword-based search queries for web search) or implicit signals (e.g., history of previously accessed artefacts in case of recommender systems), and make subjective estimates of an artefact’s relevance to the information need. Consequently, these systems are not neutral tools for lookup [274] and the choices these systems make exert systemic influence over what information is exposed and consumed at scale. These systems bear a responsibility to society to not only mitigate potential harms, like allocative and representational harms [87], but also to maximize social good [28].
@@
Representational harms may happen due to reinforcement of negative stereotypes (e.g., by disproportionately suggesting arrest record searches in ads corresponding to black-identifying first names [339] or suggesting racist stereotypes in query autocompletion [274]), by pandering to the white male gaze (e.g., by sexualizing women of color in image search results), and through erasure (e.g., underrepresenting women and other historically marginalized peoples in image search for occupational roles [206]). Allocative harms may manifest from disparate exposure in search and recommendation results [324]e.g., when women are recommended lower-paying jobs in ads [95] or by influencing traffic to websites that depend on ad-monetization. Beyond direct representational and allocative harms, these systems also hold tremendous power to shape political discourse and culture [162, 146, 168].
In light of these, there has been multiple calls [91, 276] in the IR community to study and address these potential harms. Bernard and Balog [46] report a significant rise in publications in this area after 2016, with fairness and transparency receiving the most attention. Fairness in ranking has garnered so much interests that there are numerous recent surveys [119, 402, 403, 404, 289, 103, 1, 282, 375, 228, 97] and tutorials [118, 141, 227, 123, 54] summarizing this emerging body of work, as well as shared tasks [50, 51, 51, 120] and doctoral theses (e.g., [238, 140, 331, 297]). The fairness questions have typically been framed around disparate quality-of-service [253, 392, 236, 254, 116, 70, 47, 337, 281, 391, 389, 273, 272, 115, 16, 19] and disparate exposure [394, 49, 324, 70, 325, 117, 337, 99, 393, 401, 74, 281, 268, 277, 52, 391, 332, 390, 389, 142, 272, 316, 405, 296, 173, 145, 252, 309, 255, 395, 189, 16, 56, 55, 216, 19]. We intentionally cite these individual works—and by no means is this list complete—to reemphasize the volume of papers that have been authored recently on measurement and mitigation of bias and unfairness in ranking. Consequently, several recent papers [329, 298, 299, 62, 63] have tried to systemically compare various fairness metrics proposed across these works.
@@
It is important that we critically assess how this body of research translates to real world impact. But on this, there is little in the published literature to go on. It is plausible that this research has been operationalized in, or at least has influenced the designs of, popular IR systems in recent years, but organizations who build these systems have not publicly disclosed this due to concerns of safety or for competitive reasons. Alternatively, it is also plausible that in fact many of these approaches have not been operationalized, or only been operationalized in very constrained settings, because industry adoption is lagging behind research or that existing fairness research is built on abstractions and assumptions that are incompatible with real world deployment.
@@
Studies of search logs [193, 81] have historically served an important role in IR research for understanding user and system behaviors and how the two interact. Similarly, online experimentation [211, 212, 213] has been key to validating the outputs of IR research in the real world. In contrast, to the best of our knowledge, there are very few fairness studies (e.g.[253, 254, 300]) in IR that make use of search logs, while it may be argued that the user interaction data in these logs are exactly where we should be looking to identify which social harms are common in practice and understand how exactly they manifest. Similarly, there is an urgent need for validating proposed fairness interventions from the literature through online experimentation, involving real users and real information needs, to ensure that fairness research is grounded in actual needs of the society and does not amount to just an intellectual exercise. While the lack of log-based studies and online experimentation in fairness research is likely due to the lack of access to commercially-deployed systems and corresponding log data for academic research, we must critically enquire why this manifests so much more severely in fairness research compared to other areas of IR, such as in ranking. To do so, we must expand the frame beyond questions of algorithmic fairness, and examine the very sociopolitical context in which this research is being conducted, the economic incentives and risks that shape it, and the power differentials between institutions and individuals that determine what research is allowed and who is allowed to do it [379, 381]. This lack of access to data and systems not only makes it difficult to reproduce, validate, and challenge the claims in existing fairness studies,222 For example, some studies on log-based fairness audits depend on the availability of user provided demographic attributes. In our experience, such attributes are often available only for a small subset of users—such as for signed-in users—who typically are more loyal users of the product and on average much better satisfied with system performance than the general population. Audits based on this sub-population may significantly under-report bias and unfairness issues faced by users. Better understanding of such practical challenges can motivate the community to work on specific research questions, such as how to conduct fairness audits under distributional shifts and with noisy [145, 252] or missing demographic attributes (e.g., [104]). but also limits what fairness questions the community is allowed to investigate.
@@
We should also critically reflect on what questions should or should not be framed as fairness problems, and the societal consequences of doing so. For example, one of the motivating scenarios described by Morik et al. [268] is exposure fairness for search results across different ends of the political spectrum. A similar question has also been studied by Kulshrestha et al. [217]. Casting this as a fairness issue, however, has several problematic implications and consequences. Firstly, this assumes an overly-simplistic frame in which complicated intersecting political ideologies are mapped to a linear spectrum (e.g., left vs. right and Democrats vs. Republicans) and holds the two ends static rather than a continuously shifting window of acceptable discourse [224, 147]. Furthermore, it also assumes that two ends of a political discourse have equal merit and deserve equal exposure, which amounts to algorithmic bothsidism. Finally, in arguing that builders of IR systems should shift exposure out of fairness concerns, it inadvertently normalizes the idea that it is acceptable for institutions and individuals who own these systems to exert enormous influence over public discourse of social and moral import. Instead, it is our perspective that IR needs a fundamentally different and interdisciplinary approach to these questions, one that centers on engaging and co-producing with other academic sub-fields, such as democratic theory (e.g.[94, 93, 90, 92, 171, 318, 207, 109, 110]), critical theory (e.g.[178, 179, 8, 7, 45, 44, 244, 243, 166, 176, 177, 42, 132, 124, 133, 130, 71, 111, 333, 9]), and particularly critical theory of technology (e.g.[126, 125]).
@@
We must also assess the validity of constructs that we employ in fairness research. For example, Jacobs and Wallach [188] point out that race and gender, that are often the focus of group fairness research, are contested constructs, and indeed so is the construct of fairness itself. While several papers on fairness have employed gender as a group variable [253, 117, 116, 53, 52, 316, 239, 405, 55, 215, 273, 255, 216, 56, 300], Pinney et al. [288] caution us that much more care should be taken in this practice, for example, to ensure that we respect everyone’s right to self-identify their gender and recognize the fairness concerns of non-binary peoples. Patro et al. [282] encourage us to to move beyond fairness definitions that are grounded in discrete moments and to consider the long-term impact of fairness interventions. Fairness research itself may contribute towards certain negative externalities in the long-term, such as encouraging more pervasive collection of protected demographic attributes and further intensification of data surveillance [411] of already marginalized groups in a misguided attempt to bridge the data gaps that may be responsible for the system’s disparate quality of service across groups.
Beyond fairness, there has been renewed interests in questions of transparency and explainability [406, 18], addressing misinformation [407, 218, 319, 82, 408, 311, 163], and broader ethics in IR [313].

### v4

# Search and Society: Reimagining Information Access for Radical Futures

@@ Abstract
Information retrieval (IR) research must understand and contend with the social implications of the technology it produces. Instead of adopting a reactionary strategy of trying to mitigate potential social harms from emerging technologies, the community should aim to proactively set the research agenda for the kinds of systems we should build inspired by diverse explicitly stated sociotechnical imaginaries. The sociotechnical imaginaries that underpin the design and development of information access technologies needs to be explicitly articulated, and we need to develop theories of change in context of these diverse perspectives. Our guiding future imaginaries must be informed by other academic fields, such as human-computer interaction, information sciences, media studies, design, science and technology studies, social sciences, humanities, democratic theory, and critical theory, as well as legal and policy experts, civil rights and social justice activists, and artists, among others. In this perspective paper, we motivate why the community must consider this radical shift in how we do research and what we work on, and sketch a path forward towards this transformation.
Keywords: IR and society, Sociotechnical imaginaries, Theory of change, Technology and power
@@ 1 Introduction
Information retrieval (IR) research must understand and contend with the social implications of the technology it produces. Nearly half a century ago, Belkin and Robertson (1976) concluded that IR research should acknowledge its responsibility to society and “must become both theoretically self-conscious and self-consciously based upon a social ideology”. This perspective has gained traction in the IR community in recent years. Researchers in attendance at the third Strategic Workshop in Information Retrieval in Lorne (SWIRL) (Culpepper et al., 2018) identified fairness, accountability, confidentiality, and transparency in IR (“FACT IR”) as socially consequential and strategically important research directions for the field. The following year, the FACTS-IR workshop (Olteanu et al., 2019) added “safety” as a fifth pillar. Subsequently, a large body of recent IR literature has grappled with questions of fairness, transparency, and explainability in the context of information access.
However, it is our perspective that this growing focus on fairness and ethics in IR—despite having played a critical role in bringing much-needed attention to the societal implications of IR systems and advancing the conversation about the IR community’s responsibility to broader society—operates within a severely constrained frame that leaves the many underlying values, politics, and socioeconomic incentives that guide IR research largely unchallenged. For example, faced with the applications of generative artificial intelligence (AI) for information access the IR community has focused on concerns of fair ranking and representation and limiting model “hallucinations”111 We acknowledge that the term “hallucination” anthropomorphizes AI models and its usage should be discouraged. However, given the popular usage of that term in the IR community, we make an exception here for clarity. for good reasons, but have largely ignored other significant consequences of these technologies on society, such as for the information ecosystem and how these systems concentrate power and control (Mitra et al., 2024). In machine learning (ML), there has been similar recent perspectives (Blodgett et al., 2020; Miceli et al., 2022) that, for example, calls for shifting the lens from fairness and bias to the power differentials that exists between those who build technology, those who use it, and those who are subject to it. Others in the ML community have brought attention to the questions of how these technologies shift power (Kalluri et al., 2020) and simultaneously constrain ethics interventions in practice (Widder et al., 2023) and shape our collective sociotechnical futures. Even by accepting the frame that we should develop fairness and transparency mechanisms for certain systems, we may inadvertently ignore the alternative perspective that some of these technologies should be dismantled, not made fairer, nor more transparent (Barocas et al., 2020; Wilkinson et al., 2023; Merchant, 2023). Consequently, at the recent fourth SWIRL workshop [^2] https://sites.google.com/view/swirl2025, researchers in attendance called for expanding the “FACTS-IR” framing to center IR research on societal, democratic, and emancipatory values. Similar sentiments, including re-centering IR on societal needs and informing IR research with democratic theories, were also discussed at the first Search Futures workshop (Azzopardi et al., 2024b).
What do we propose?
@@
In this paper, we argue that IR research, instead of adopting a reactionary strategy of trying to mitigate potential social harms from emerging technologies by developing new fairness and transparency interventions, should aim to proactively set the research agenda for the kinds of systems we should build inspired by diverse explicitly stated sociotechnical imaginaries. Towards that goal, IR research needs to explicitly articulate the sociotechnical imaginaries (Jasanoff and Kim, 2009, 2015) that underpin the design and development of information access technologies, and develop theories of change (Weiss, 1995; Brest, 2010; Taplin and Clark, 2012; Wikipedia contributors, 2013). Jasanoff and Kim (2015) define sociotechnical imaginaries as “collectively held, institutionally stabilized, and publicly performed visions of desirable futures, animated by shared understandings of forms of social life and social order attainable through, and supportive of, advances in science and technology”. These shared visions do not only imagine but also co-produce our futures through development and government of digital technologies (Mager and Katzenbach, 2021). Diverse imaginaries promoted by different corporations, professional communities, political organizations, and social movements can coexist “in tension or in a productive dialectical relationship” (Jasanoff and Kim, 2015).
@@
In technological research and development, these diverse, and often diverging, perspectives and visions that guide the community frequently remain implicit and unstated (Wilkinson et al., 2023) despite the significant influence they exert on what the community focuses on and produces. Because of the consequential role that access to information plays in political participation by citizens in democratic societies and social transformation (Higgins and Gregory, 2013; Polizzi, 2020; Goldstein, 2020; Correia, 2002; González, 2021), and as a social determinant of health (Moretti et al., 2012) and economic progress (Yu, 2002; Mutula, 2008), it is even more important to critically reflect on the values and motivations that guide the design and deployment of popular information access systems. What sociotechnical futures do IR researchers and system designers envision and how do they influence the design of current and future IR systems? Whose sociotechnical imaginaries are granted normative status and what myriad of radically alternative futures are we overlooking? For example, what are the implications of the reliance of popular search and social media platforms on advertising as the primary source of revenue generation (Ang, 2022) and how does Big Tech’s (Oremus, 2017) increasing dominance over academic research (Whittaker, 2021) influences and/or homogenizes the kinds of IR systems we build? What is our role, as IR researchers, to safeguard communities from falling victim to crisis of imaginations (Haiven, 2014) and how do we become more open to welcoming influences from radically new sociotechnical imaginaries? For example, what would IR systems look like if designed for futures informed by feminist, queer, decolonial, anti-racist, anti-casteist, anti-ableist, and abolitionist thoughts, and if the focus of IR research was not to prop up colonial cisheteropatriarchal capitalist structures but to dismantle them? We believe that explicating, critiquing, and consciously choosing the values and sociotechnical imaginaries that shape IR research is critical to realizing positive social outcomes through IR research. As Benjamin (2024) argues, exercising our imagination is “an invitation to rid our mental and social structures from the tyranny of dominant imaginaries”, or as Le Guin put it:
@@
The Wave in the Mind: Talks and Essays on
the Writer, the Reader, and the Imagination
 (Le Guin, 2004)
@@
For IR to concretely support diverse sociotechnical imaginaries the research community also needs to develop their own theories of change. Theory of change (Weiss, 1995; Brest, 2010; Taplin and Clark, 2012; Wikipedia contributors, 2013) can be defined as a participatory process whereby groups and stakeholders articulate their long-term goals and identify necessary preconditions in a planning process. Consequently, in IR adjacent fields—e.g., in the Fairness, Accountability, and Transparency (FAccT) [^3] https://facctconference.org community—there has been similar recent calls to reimagine our sociotechnical futures (da Hora et al., 2024) and develop theories of change (Wilkinson et al., 2023) to make explicit the visions for desired futures of responsible computing and the strategic pathways that lead to those desired futures. Here, we argue that IR research similarly needs to both explicitly articulate and support diverse imagined futures and develop corresponding theories of change for how new information access technologies can take us towards these desired worlds. Theories of change in this context may benefit IR research by encouraging community members to make their goals and assumptions explicit, making it feasible to test stated theories, encouraging the community to work towards building consensus, and even aid in developing potential means of evaluation of desired progress (Weiss, 1995).
@@
Maximizing social good and minimizing harm in this context should not just be the concerns of the few in our community working on fairness, transparency, ethics, and related areas, but the domain of all IR research that should be guided by theories of change towards these envisioned futures. In this context, we largely agree with the perspective of Belkin and Robertson (1976) but diverge on a critical point which is that we believe IR research should not seek any singular notions of “universal” social ideology but explicitly adopt pluralistic humanistic and emancipatory values and make space for diverse visions and perspectives.
@@
Consequently, the task for IR researchers here is not to put themselves in positions to pick the guiding social ideologies nor push technosolutionism to address today’s social problems. Instead, our guiding future imaginaries must be co-developed with scholars from diverse fields such as human-computer interaction (HCI), information sciences, media studies, design, science and technology studies (STS), social sciences, and political sciences, as well as legal and policy experts, civil rights and social justice activists, and artist, to name a few. Not all sociotechnical imaginaries are equal in this respect, and what futures guide our research must be informed by the values and ethics of our community, which should be constantly discussed, debated, and challenged by the community as part of our sociotechnical research and be open to external critique. To summarize, as a research community we should invest our energies and resources to: (i) Nurture digital spaces where radical visions and projects for human emancipation, social progress, and equity and justice can take shape, (ii) encourage experimentation within our research community with new approaches to information access informed by new sociotechnical imaginaries cross-pollinating through interdisciplinary scholarship, and (iii) ensure that the tools and artefacts we produce as a community do not uphold systems of oppression nor contribute towards other systemic social harms.
@@
So far, we have argued that IR research should explicate the sociotechnical futures we want to realize and develop theories of change towards these desired futures. In Section 2 we present a brief overview of existing literature on fairness and ethics in IR, and share our critical perspectives on it to motivate our work. Then, in Section 3 we draw from relevant movements in IR-adjacent fields, specifically those with explicit values and prefigurative politics; we share some ideas on how the community can get started on this journey; and who should be doing this work. In Section 4 we motivate why now is an appropriate time for the community to consider this shift. We conclude in Section 5 with some final remarks on potential pitfalls and desired outcomes. Our goal with this paper is to raise sociopolitical consciousness in the IR research community so that we all see our research embedded in projects of future world making and recognize our collective responsibility to affect social good; and to dismantle the artificial separation between the work on fairness and ethics in IR and the rest of IR research.
@@ 2 Background
IR systems act as intermediaries between information seekers and information artefacts. These artefacts may represent: (i) economic and other opportunities for consumers, (ii) monetization opportunities for content creators and publishers, (iii) specific sociopolitical frames and ideologies, and (iv) lenses to view individuals and groups that are subjects of representation by the content. These systems infer the information need from highly incomplete expressions of interests (e.g., short keyword-based search queries for web search) or implicit signals (e.g., history of previously accessed artefacts in case of recommender systems), and make subjective estimates of an artefact’s relevance to the information need. Consequently, these systems are not neutral tools for lookup (Noble, 2018) and the choices these systems make exert systemic influence over what information is exposed and consumed at scale. These systems bear a responsibility to society to not only mitigate potential harms, like allocative and representational harms (Crawford, 2017), but also to maximize social good.
@@
Representational harms may happen due to reinforcement of negative stereotypes (e.g., by disproportionately suggesting arrest record searches in ads corresponding to searches for black-identifying first names (Sweeney, 2013) or suggesting racist stereotypes in query autocompletion (Noble, 2018)), by pandering to the white male gaze (e.g., by sexualizing women of color in search results (Noble, 2018; Urman and Makhortykh, 2024)), and through erasure (e.g., underrepresenting women and other historically marginalized peoples in image search for occupational roles (Kay et al., 2015)). Allocative harms may manifest from disparate exposure in search and recommendation results (Singh and Joachims, 2018)e.g., when women are recommended lower-paying jobs in ads (Datta et al., 2014) or by influencing traffic to websites that depend on ad-monetization. Beyond direct representational and allocative harms, these systems also hold tremendous power to shape political discourse and culture (Grimmelmann, 2008; Gillespie, 2019; Hallinan and Striphas, 2016).
In light of these, there has been multiple calls (Culpepper et al., 2018; Olteanu et al., 2019) in the IR community to study and address these potential harms. Bernard and Balog (2023) report a significant rise in publications in this area after 2016, with fairness and transparency receiving the most attention. The increasing focus on these sociotechnical aspects of information access has been at least partly in response to recent advances in foundation models (Bommasani et al., 2021) and their implications for the future of information access (Shah and Bender, 2022; Mitra et al., 2024).
Fairness in ranking has garnered so much interests that there are numerous recent surveys (Ekstrand et al., 2021; Zehlike et al., 2021, 2022a, 2022b; Pitoura et al., 2022; Dinnissen and Bauer, 2022; Aalam et al., 2022; Patro et al., 2022; Wang et al., 2023; Li et al., 2023; Deldjoo et al., 2023) and tutorials (Ekstrand et al., 2019; Gao and Shah, 2020; Li et al., 2021; Fang et al., 2022; Bigdeli et al., 2022) summarizing this emerging body of work, as well as shared tasks (Biega et al., 2020, 2021; Ekstrand et al., 2023). The fairness questions have typically been framed around disparate quality-of-service—e.g.(Mehrotra et al., 2017, 2018; Neophytou et al., 2022; Wu et al., 2024)—and disparate exposure—e.g.(Biega et al., 2018; Singh and Joachims, 2018, 2019; Diaz et al., 2020; Zehlike and Castillo, 2020; Patro et al., 2020; Wu et al., 2022b). Several recent works (Smith and Beattie, 2022; Raj and Ekstrand, 2020, 2022; Boratto et al., 2022, 2023) have also systemically compared various fairness metrics proposed in the literature.
Beyond fairness, there has been renewed interests in questions of transparency and explainability (Zhang et al., 2020; Anand et al., 2022), addressing misinformation (Zhou and Zafarani, 2018; Kumar and Shah, 2018; Sharma et al., 2019; Collins et al., 2021; Zhou and Zafarani, 2020; Saracco et al., 2021; Guo et al., 2022), and broader ethical concerns in IR (Schedl et al., 2022). Transparency in IR covers a broad range of scenarios and concerns. Examples include transparency about how the system behaves (Singh and Anand, 2019; Verma and Ganguly, 2019; Singh and Anand, 2018; Zhuang et al., 2020) and how data subjects are represented in search results (Biega et al., 2017; Li et al., 2022). Transparency needs may specifically arise in the context of how information and knowledge access systems modulate what and who get exposure and influence how we see ourselves and others (Cortiñas-Lorenzo et al., 2024). Different notions of transparency may be relevant here, including but not limited to: System transparency (i.e., how does the system work?), procedural transparency (i.e., in what social norms and processes is the system use embedded?), or transparency of outcomes (i.e., what is the impact of the system’s use on individuals and society?). New transparency needs (Liao and Vaughan, 2023) may also arise in the context of emerging technologies, such as large language models (LLMs) (OpenAI, 2023; Thoppilan et al., 2022; Chowdhery et al., 2022; Touvron et al., 2023).
2.1 Critical perspectives on fairness and ethics research in IR
@@
It is important that we critically assess how the body of fairness and ethics research in IR translates to real world impact. But on this, there is little in the published literature to go on. It is plausible, for example, that the research on ranking fairness has been operationalized in, or at least has influenced the designs of, popular IR systems in recent years, but institutions who build these systems have rarely publicly disclosed any information in that regard, may be to obfuscate details of system design from bad actors or for competitive reasons. Alternatively, it is also plausible that in fact many of these approaches have not been operationalized, or only been operationalized in very constrained settings, because industry adoption is lagging behind research or that existing fairness research is built on abstractions and assumptions that are incompatible with real world deployment.
@@
Studies of search logs (Jiang et al., 2013; Chuklin et al., 2015) have historically served an important role in IR research for understanding user and system behaviors and how the two interact. Similarly, online experimentation (Kohavi et al., 2007, 2009, 2020) has been key to validating the outputs of IR research in the real world. In contrast, to the best of our knowledge, there are very few fairness studies—e.g.(Mehrotra et al., 2017, 2018; Raj et al., 2023)—in IR that make use of search logs, while it may be argued that the user interaction data in these logs are exactly where we should be looking to identify which social harms are common in practice and understand how exactly they manifest. Similarly, there is an urgent need for validating proposed fairness interventions from the literature through online experimentation, involving real users and real information needs, to ensure that fairness research is grounded in actual needs of the society and does not amount to just academic intellectual exercises. While the lack of log-based studies and online experimentation in fairness research is likely due to the lack of access to commercially-deployed systems and corresponding log data for academic research, we must critically enquire why this manifests so much more severely in fairness research compared to other areas of IR, such as in ranking. To do so, we must expand the frame beyond questions of algorithmic fairness, and examine the very sociopolitical context in which this research is being conducted, the economic incentives and risks that shape it, and the power differentials between institutions and individuals that determine what research is allowed and who is allowed to do it (Whittaker, 2021; Widder et al., 2023). This lack of access to data and systems not only makes it difficult to reproduce, validate, and challenge the claims in existing fairness studies,444 For example, some studies on log-based fairness audits depend on the availability of user provided demographic attributes. In our experience, such attributes are often available only for a small subset of users—such as for signed-in users—who typically are more loyal users of the product and on average much better satisfied with system performance than the general population. Audits based on this sub-population may significantly under-report bias and unfairness issues faced by users. Better understanding of such practical challenges can motivate the community to work on specific research questions, such as fairness considerations under distributional shifts and with noisy (Ghazimatin et al., 2022; Mehrotra and Vishnoi, 2022) or missing demographic attributes (Lazovich et al., 2022; Do and Usunier, 2022). but also limits what fairness questions the community is allowed to investigate.
@@
We should also critically reflect on what questions should or should not be framed as fairness problems, and the societal consequences of doing so. For example, one of the motivating scenarios described by Morik et al. (2020) is exposure fairness for search results across different ends of the political spectrum. A similar question has also been studied by Kulshrestha et al. (2017). Casting this as a fairness issue, however, has several problematic implications and consequences. Firstly, this assumes an overly-simplistic frame in which complicated intersecting political ideologies are mapped to a linear spectrum (e.g., left vs. right and Democrats vs. Republicans) and holds the two ends static rather than a continuously shifting window of acceptable discourse (Lehman, 2014; Giridharadas, 2019a). Furthermore, it also assumes that two ends of a political discourse have equal merit and deserve equal exposure, which amounts to algorithmic bothsidism. Finally, in arguing that builders of IR systems should shift exposure out of fairness concerns, it inadvertently normalizes the idea that it is acceptable for institutions and individuals who own these systems to exert enormous influence over public discourse of social and moral import. Instead, it is our perspective that IR needs a fundamentally different and cross-disciplinary approach to these questions, one that centers on engaging and co-producing with other academic sub-fields, such as STS (Hackett et al., 2008) and critical theory of technology (Feenberg et al., 1991; Feenberg, 2002).
@@
We must also assess the validity of constructs that we employ in fairness research. For example, Jacobs and Wallach (2021) point out that race and gender, that are often the focus of group fairness research, are contested constructs, and indeed so is the construct of fairness itself. While several papers on fairness have employed gender as a group variable, Pinney et al. (2023) caution us that much more care should be taken in this practice, for example, to ensure that we respect everyone’s right to self-identify their gender and recognize the fairness concerns of non-binary peoples. Patro et al. (2022) encourage us to to move beyond fairness definitions that are grounded in discrete moments and to consider the long-term impact of fairness interventions. Fairness research itself may contribute towards certain negative externalities in the long-term, such as encouraging more pervasive collection of protected demographic attributes and further intensification of data surveillance (Zuboff, 2023) of already marginalized groups in a misguided attempt to bridge the data gaps that may be responsible for the system’s disparate quality of service across groups.